# On the Source Model Key Agreement Problem

### v2

@@ B.3 Proof of Theorem 3
Note that when $P(Y=2)=0$, i.e., when $Y\in\{0,1\}$, the random variable $Z$ can be expressed as $Z=X\oplus Y$. In this case, proof of Theorem 2 shows that one can find a random variable $T$ such that $I(Z;T|X,Y)=I(Z;T)=I(X;Y|T)=0$. Lemma 1 would then imply any $P_{X,Y}$ satisfying $P_{Y}(2)=0$ belongs to the convex hull of the set $\mathcal{A}(P_{Z})$ as defined in (12). More specifically, let $\alpha=P_{Z}(0)$ and $\bar{\alpha}=P_{Z}(1)=1-\alpha$ and let us denote $\mathcal{A}(P_{Z})$ by $\mathcal{A}(\alpha)$. Then, for any $r,t\in[0,1]$, the following joint distribution belongs to the convex hull of the set $\mathcal{A}(\alpha)$:
@@
Observe that when $P(Y=0)=0$, the mapping from $X$ and $Y\in\{1,2\}$ to $Z$ is again an XOR map. A similar argument shows that then imply any $P_{X,Y}$ satisfying $P_{Y}(0)=0$ belongs to the convex hull of the set $\mathcal{A}(\alpha)$. In other words, for any $s,\ell\in[0,1]$ the following joint distribution belongs to the convex hull of the set $\mathcal{A}(\alpha)$:
@@
Observe that for any $\omega\in[0,1]$ the linear combination
@@
$r=\min\left\{1,\frac{\bar{\alpha}p_{00}}{\alpha p_{10}}\right\}$
$\omega=\frac{p_{00}}{\alpha r},~{}t=\frac{\alpha p_{10}}{\bar{\alpha}p_{00}}r,~{}s=\frac{\alpha p_{12}r}{\bar{\alpha}(\alpha r-p_{00})},~{}\ell=\frac{p_{02}r}{\alpha r-p_{00}}.$
@@
$\frac{p_{00}}{\alpha}+\frac{p_{12}}{\bar{\alpha}}\leq 1,~{}\frac{p_{02}}{\alpha}+\frac{p_{10}}{\bar{\alpha}}\leq 1.$
@@
Consequently, one can identify a random variable $T$ such that $I(Z;T|X,Y)=I(Z;T)=I(X;Y|T)=0$ holds for $P_{X,Y}$ provided that the conditions
$\frac{p_{00}}{\alpha}+\frac{p_{12}}{\bar{\alpha}}\leq 1,~{}\frac{p_{02}}{\alpha}+\frac{p_{10}}{\bar{\alpha}}\leq 1$
@@ B.4 Proof of Theorem 4
Let the alphabets of $X$ and $Y$ take values in $\{0,1,2,\cdots,s_{1}-1\}$ and $\{0,1,2,\cdots,s_{2}-1\}$ respectively. Let $Z=f(X,Y)$. The function $f$ corresponds to a table
A table is considered valid if the equality $\Psi(X;Y|Z)=I(X;Y)$ holds for all distributions $P_{X,Y}$. Since any distribution over a subset of $\mathcal{X}\times\mathcal{Y}$ can be extended to a distribution over $\mathcal{X}\times\mathcal{Y}$, it follows that if a table is valid, all of its sub-tables will also be valid. Consider a $2\times 2$ table. As stated in Theorem 2, the equality $\Psi(X;Y|Z)=I(X;Y)$ does not hold for any arbitrary distribution $P_{X,Y}$ when the functions are AND, OR, or $Z_{1}$. It is also evident that if $Z=f(X,Y)$ is a one-to-one map, $\Psi(X;Y|Z)\leq I(X;Y|Z)=0$ and the equality $\Psi(X;Y|Z)=I(X;Y)$ would not hold for every $P_{X,Y}$. Consequently, for $2\times 2$ tables, the following combinations are invalid if $A,B,C$ and $D$ are distinct symbols:
@@
$T(i_{1},j_{1})=T(i_{2},j_{2})\neq T(i_{2},j_{1})=T(i_{1},j_{2})$
In both cases $T(i_{1},j_{1})=T(i_{2},j_{2})$ for every $i_{1}\neq i_{2}$ and $j_{1}\neq j_{2}$. In other words, if we move from one cell to another cell by changing both the row and the column, the value of the new cell must be equal to the value of the old cell. Note that in a $s_{1}\times s_{2}$ table where $(s_{1},s_{2})\neq(2,2)$, we can move from a cell to all the other cells by a sequence of such moves. For instance, in a $2\times 3$ table, we can move from $(1,1)$ to $(1,2)$ by the following sequence:
$(1,1)\rightarrow(2,2)\rightarrow(3,1)\rightarrow(1,2).$
@@ B.5 Proof of Theorem 5
Let $P_{J|X^{\prime}Y^{\prime}Z^{\prime}}$ be a minimizer for the optimization problem in $\hat{\Psi}(X^{\prime},Y^{\prime}\|Z^{\prime})$, i.e.,
@@
We know that such a minimizer exists because of the cardinality bound on $J$ is available for $\hat{\Psi}(X^{\prime},Y^{\prime}\|Z^{\prime})$.
We define the joint distribution between $J$ and $(X,Y,Z,X^{\prime},Y^{\prime},Z^{\prime})$ as $P_{J|X^{\prime}Y^{\prime}Z^{\prime}}P_{X^{\prime}Y^{\prime}Z^{\prime}XYZ}$. This construction implies the Markov chain
@@
$\leq I(XY;Z^{\prime}|Z)+I(Y^{\prime}Z^{\prime};X|X^{\prime})+I(X^{\prime}Z^{\prime};Y|Y^{\prime})+I(X;Y|X^{\prime}Y^{\prime}Z^{\prime}).$ (76)
@@
$I(X;Y|J)+I(XY;J|Z)-I(XX^{\prime};YY^{\prime}|J)-I(XX^{\prime}YY^{\prime};J|Z^{\prime})$ $\leq I(XY;Z^{\prime}|Z),$ (77)
$I(XX^{\prime};YY^{\prime}|J)+I(XX^{\prime}YY^{\prime};J|Z^{\prime})-I(XX^{\prime};Y^{\prime}|J)-I(XX^{\prime}Y^{\prime};J|Z^{\prime})$ $\leq I(XX^{\prime}Z^{\prime};Y|Y^{\prime}),$ (78)
$I(XX^{\prime};Y^{\prime}|J)+I(XX^{\prime}Y^{\prime};J|Z^{\prime})-I(X^{\prime};Y^{\prime}|J)-I(X^{\prime}Y^{\prime};J|Z^{\prime})$ $\leq I(Y^{\prime}Z^{\prime};X|X^{\prime}).$ (79)
@@
$I(X;Y|J)+I(XY;J|Z)-I(XX^{\prime};YY^{\prime}|J)-I(XX^{\prime}YY^{\prime};J|Z^{\prime})$
$\leq I(XY;JZ^{\prime}|Z)-I(XX^{\prime}YY^{\prime};J|Z^{\prime})$
$\leq I(XY;Z^{\prime}|Z)+I(XYZ;J|Z^{\prime})-I(XX^{\prime}YY^{\prime};J|Z^{\prime})$
$\leq I(XY;Z^{\prime}|Z)$
@@
$I(XX^{\prime};YY^{\prime}|J)+I(XX^{\prime}YY^{\prime};J|Z^{\prime})-I(XX^{\prime};Y^{\prime}|J)-I(XX^{\prime}Y^{\prime};J|Z^{\prime})$
$=I(XX^{\prime};Y|JY^{\prime})+I(Y;J|XX^{\prime}Y^{\prime}Z^{\prime})$
$=I(XX^{\prime};Y|JY^{\prime})-I(Y;J|XX^{\prime}Y^{\prime}Z^{\prime})$ (80)
$\leq I(XX^{\prime}Z^{\prime};Y|Y^{\prime})-I(J;Y|Y^{\prime})$
$\leq I(XX^{\prime}Z^{\prime};Y|Y^{\prime})$
@@
Finally, the inequality (79) similarly follows from
$I(XX^{\prime};Y^{\prime}|J)+I(XX^{\prime}Y^{\prime};J|Z^{\prime})-I(X^{\prime};Y^{\prime}|J)-I(X^{\prime}Y^{\prime};J|Z^{\prime})$ (81)
$=I(X;Y^{\prime}|X^{\prime}J)+I(X;J|X^{\prime}Y^{\prime}Z^{\prime})$ (82)
$=I(X;Y^{\prime}|X^{\prime}J)-I(X;J|X^{\prime}Y^{\prime}Z^{\prime})$ (83)
$\leq I(Y^{\prime}Z^{\prime};X|X^{\prime})-I(J;X|X^{\prime})$ (84)
$\leq I(Y^{\prime}Z^{\prime};X|X^{\prime})$ (85)
@@ B.6 Proof of Theorem 6
Let $F_{1:r}=(F_{1},\cdots,F_{r})$. Take some arbitrary $P_{J|X,Y,Z}$ and suppose $(X^{n},Y^{n},Z^{n},J^{n})$ are $n$ i.i.d. copies according to $P_{X,Y,Z}P_{J|X,Y,Z}$. Define the auxiliary random variables as $V_{i}=(J^{i-1},Z_{i+1}^{n},F_{1:r})$ and $U_{i}=K_{A}$. The following chains of equations follows:
@@
=$\overset{(b)}{=}n\epsilon_{n}+\sum_{i=1}^{n}I(K_{A};J_{i}|J^{i-1},Z_{i+1}^{n},F)-I(K_{A};Z_{i}|J^{i-1},Z_{i+1}^{n},F)+H(K_{A}|J^{n},F_{1:r})$
=$\overset{(c)}{=}n\epsilon_{n}+n\delta_{n}+\sum_{i=1}^{n}I(U_{i};J_{i}|V_{i})-I(U_{i};Z_{i}|V_{i})+I(K_{A};K_{B}|J^{n},F_{1:r})$
@@
where $\epsilon_{n},\delta_{n}\rightarrow 0$ as $n\rightarrow\infty$; steps (a) and (c) are derived using Fano’s inequality, while (b) follows from the Marton-Körner identity, and (d) is a consequence of the Data Processing Inequality (DPI). To establish (e), we proceed as follows:
$I(X^{n};Y^{n}|J^{n})$ =$\overset{(a)}{=}I(X^{n},F_{1};Y^{n}|J^{n})\overset{(b)}{\geq}I(X^{n};Y^{n}|J^{n},F_{1})\overset{(c)}{=}I(X^{n};Y^{n},F_{2}|J^{n},F_{1})\overset{(d)}{\geq}I(X^{n};Y^{n}|J^{n},F_{1},F_{2})$
@@
where $(a)$ and $(b)$ follow from Markov chains $(Y^{n},F^{i-1})-\!\!\!\!\minuso\!\!\!\!-X^{n}-\!\!\!\!\minuso\!\!\!\!-F_{i}$ for odd $i$ and $(X^{n},F^{i-1})-\!\!\!\!\minuso\!\!\!\!-Y^{n}-\!\!\!\!\minuso\!\!\!\!-F_{i}$ for even $i$ respectively. Also $(b)$ and $(d)$ are derived from chain rule.
Let $Q$ be uniformly distributed over $\{1,2,\cdots,n\}$ and independent of $X,Y,Z,J,F_{1:r}$. Define the random variables $V=(V_{Q},Q),U=U_{Q},Z=Z_{Q}$ and $J=J_{Q}$. Under these definitions, inequality (87) simplifies to:
@@
Additionally, the auxiliary random variables $V_{i}=(J^{i-1},Z_{i+1}^{n},F_{1:r})$ and $U_{i}=K_{A}$ imply the Markov chain relationship $(U,V)-\!\!\!\!\minuso\!\!\!\!-(X,Y)-\!\!\!\!\minuso\!\!\!\!-(J,Z)$. Assuming that $n$ is sufficiently large, to prove constraint (24) for the random variables $V_{i},J_{i},Z_{i}$ and $F_{1:r}$ as defined above, we proceed as follows:
@@
=$\overset{(a)}{=}\frac{1}{n}\sum_{i=1}^{n}I(F_{1:r};Z_{i}|Z_{i+1}^{n})+\frac{1}{n}\sum_{i=1}^{n}I(Z_{i+1}^{n};J_{i}|J^{i-1},F_{1:r})$
$\overset{(b)}{\leq}\frac{1}{n}I(Z^{n};F_{1:r})+\frac{1}{n}\sum_{i=1}^{n}I(J^{i-1},Z_{i+1}^{n},F;J_{i})$
@@
where step (a) follows from Marton-Körner identity, (b) is derived using the chain rule, and (c) utilizes the assumption
@@
Proof of the cardinality bounds are standard and we skip them. The second part of this theorem for the special case of $Z=f(X,Y)$ follows from Lemma 3 (after applying the proper change of variables). ∎
@@
Suppose that $Z=g(X)$. For every joint distribution $P_{V,X,Y}$, the maximizer of $I(U;Y|V)-I(U;Z|V)$ over all $P_{U|V,X}$ is achieved when $H(X|U,V,Z)=0$.

### v3

@@ B.3 Proof of Theorem 3
Note that when $P(Y=2)=0$, i.e., when $Y\in\{0,1\}$, the random variable $Z$ can be expressed as $Z=X\oplus Y$. In this case, the proof of Theorem 2 shows that one can find a random variable $T$ such that $I(Z;T|X,Y)=I(Z;T)=I(X;Y|T)=0$. Lemma 1 would then imply that any $P_{X,Y}$ satisfying $P_{Y}(2)=0$ belongs to the convex hull of the set $\mathcal{A}(P_{Z})$ as defined in (18). More specifically, let $\alpha=P_{Z}(0)$ and $\bar{\alpha}=P_{Z}(1)=1-\alpha$, and let us denote $\mathcal{A}(P_{Z})$ by $\mathcal{A}(\alpha)$. Then, for any $r,t\in[0,1]$, the following joint distribution belongs to the convex hull of the set $\mathcal{A}(\alpha)$:
@@
In the above table, the values of $X$ are listed in the first column (i.e., $X\in\{0,1\}$) and the values of $Y$ are listed in the first row (i.e., $Y\in\{0,1,2\}$). The remaining cells of the table contain probabilities.
Observe that when $P(Y=0)=0$, the mapping from $X$ and $Y\in\{1,2\}$ to $Z$ is again an XOR map. A similar argument then shows that any $P_{X,Y}$ satisfying $P_{Y}(0)=0$ belongs to the convex hull of the set $\mathcal{A}(\alpha)$. In other words, for any $s,\ell\in[0,1]$, the following joint distribution belongs to the convex hull of the set $\mathcal{A}(\alpha)$:
@@
Observe that for any $\omega\in[0,1]$, the linear combination
@@
$r=\min\left\{1,\frac{\bar{\alpha}p_{00}}{\alpha p_{10}}\right\},$
$\omega=\frac{p_{00}}{\alpha r},\quad t=\frac{\alpha p_{10}}{\bar{\alpha}p_{00}}r,\quad s=\frac{\alpha p_{12}r}{\bar{\alpha}(\alpha r-p_{00})},\quad\ell=\frac{p_{02}r}{\alpha r-p_{00}}.$
@@
$\frac{p_{00}}{\alpha}+\frac{p_{12}}{\bar{\alpha}}\leq 1,\quad\frac{p_{02}}{\alpha}+\frac{p_{10}}{\bar{\alpha}}\leq 1$
@@
Consequently, one can identify a random variable $T$ such that $I(Z;T|X,Y)=I(Z;T)=I(X;Y|T)=0$ holds for $P_{X,Y}$, provided that the conditions
$\frac{p_{00}}{\alpha}+\frac{p_{12}}{\bar{\alpha}}\leq 1,\quad\frac{p_{02}}{\alpha}+\frac{p_{10}}{\bar{\alpha}}\leq 1$
@@ B.4 Proof of Theorem 4
Let the alphabets of $X$ and $Y$ take values in $\{0,1,2,\cdots,s_{1}-1\}$ and $\{0,1,2,\cdots,s_{2}-1\}$, respectively. Let $Z=f(X,Y)$. The function $f$ corresponds to a table
In the above table, the values of $X$ are listed in the first column (i.e., $X\in\{0,1,\dots\}$) and the values of $Y$ are listed in the first row (i.e., $Y\in\{0,1,\dots\}$). The remaining cells of the table contain the values $f(x,y)$. For example, the entry in the second row and second column is $f(0,0)=f(X=0,Y=0)$.
A table is considered valid if the equality $\Psi(X;Y|Z)=I(X;Y)$ holds for all distributions $P_{X,Y}$. Since any distribution over a subset of $\mathcal{X}\times\mathcal{Y}$ can be extended to a distribution over $\mathcal{X}\times\mathcal{Y}$, it follows that if a table is valid, all of its sub-tables will also be valid. Consider a $2\times 2$ table. As stated in Theorem 2, the equality $\Psi(X;Y|Z)=I(X;Y)$ does not hold for any arbitrary distribution $P_{X,Y}$ when the functions are AND, OR, or $Z_{1}$. It is also evident that if $Z=f(X,Y)$ is a one-to-one map, then $\Psi(X;Y|Z)\leq I(X;Y|Z)=0$, and the equality $\Psi(X;Y|Z)=I(X;Y)$ would not hold for every $P_{X,Y}$. Consequently, for $2\times 2$ tables, the following combinations are invalid if $A,B,C$, and $D$ are distinct symbols:
@@
$T(i_{1},j_{1})=T(i_{2},j_{2})\neq T(i_{2},j_{1})=T(i_{1},j_{2}).$
In both cases, $T(i_{1},j_{1})=T(i_{2},j_{2})$ for every $i_{1}\neq i_{2}$ and $j_{1}\neq j_{2}$. In other words, if we move from one cell to another cell by changing both the row and the column, the value of the new cell must be equal to the value of the old cell. Note that in an $s_{1}\times s_{2}$ table where $(s_{1},s_{2})\neq(2,2)$, we can move from a cell to all the other cells by a sequence of such moves. For instance, in a $2\times 3$ table, we can move from $(1,1)$ to $(1,2)$ by the following sequence:
$(1,1)\rightarrow(2,2)\rightarrow(1,3)\rightarrow(1,2).$
@@ B.5 Proof of Theorem 5
Let $P_{J|X^{\prime},Y^{\prime},Z^{\prime}}$ be a minimizer for the optimization problem in $\hat{\Psi}(X^{\prime},Y^{\prime}\|Z^{\prime})$, i.e.,
@@
We know that such a minimizer exists because a cardinality bound on $J$ is available for $\hat{\Psi}(X^{\prime},Y^{\prime}\|Z^{\prime})$.
We define the joint distribution between $J$ and $(X,Y,Z,X^{\prime},Y^{\prime},Z^{\prime})$ as $P_{J|X^{\prime},Y^{\prime},Z^{\prime}}P_{X^{\prime},Y^{\prime},Z^{\prime},X,Y,Z}$. This construction implies the Markov chain
@@
$\leq I(X,Y;Z^{\prime}|Z)+I(Y^{\prime},Z^{\prime};X|X^{\prime})+I(X^{\prime},Z^{\prime};Y|Y^{\prime})+I(X;Y|X^{\prime},Y^{\prime},Z^{\prime}).$ (83)
@@
$I(X;Y|J)+I(X,Y;J|Z)-I(X,X^{\prime};Y,Y^{\prime}|J)-I(X,X^{\prime},Y,Y^{\prime};J|Z^{\prime})$ $\leq I(X,Y;Z^{\prime}|Z),$ (84)
$I(X,X^{\prime};Y,Y^{\prime}|J)+I(X,X^{\prime},Y,Y^{\prime};J|Z^{\prime})-I(X,X^{\prime};Y^{\prime}|J)-I(X,X^{\prime},Y^{\prime};J|Z^{\prime})$ $\leq I(X,X^{\prime},Z^{\prime};Y|Y^{\prime}),$ (85)
$I(X,X^{\prime};Y^{\prime}|J)+I(X,X^{\prime},Y^{\prime};J|Z^{\prime})-I(X^{\prime};Y^{\prime}|J)-I(X^{\prime},Y^{\prime};J|Z^{\prime})$ $\leq I(Y^{\prime},Z^{\prime};X|X^{\prime}).$ (86)
@@
$I(X;Y|J)+I(X,Y;J|Z)-I(X,X^{\prime};Y,Y^{\prime}|J)-I(X,X^{\prime},Y,Y^{\prime};J|Z^{\prime})$
$\leq I(X,Y;J,Z^{\prime}|Z)-I(X,X^{\prime},Y,Y^{\prime};J|Z^{\prime})$
$\leq I(X,Y;Z^{\prime}|Z)+I(X,Y,Z;J|Z^{\prime})-I(X,X^{\prime},Y,Y^{\prime};J|Z^{\prime})$
$\leq I(X,Y;Z^{\prime}|Z),$
@@
$I(X,X^{\prime};Y,Y^{\prime}|J)+I(X,X^{\prime},Y,Y^{\prime};J|Z^{\prime})-I(X,X^{\prime};Y^{\prime}|J)-I(X,X^{\prime},Y^{\prime};J|Z^{\prime})$
$=I(X,X^{\prime};Y|J,Y^{\prime})+I(Y;J|X,X^{\prime},Y^{\prime},Z^{\prime})$
$=I(X,X^{\prime};Y|J,Y^{\prime})-I(Y;J|X,X^{\prime},Y^{\prime},Z^{\prime})$ (87)
$\leq I(X,X^{\prime},Z^{\prime};Y|Y^{\prime})-I(J;Y|Y^{\prime})$
$\leq I(X,X^{\prime},Z^{\prime};Y|Y^{\prime}),$
@@
Finally, the inequality (86) follows similarly from
$I(X,X^{\prime};Y^{\prime}|J)+I(X,X^{\prime},Y^{\prime};J|Z^{\prime})-I(X^{\prime};Y^{\prime}|J)-I(X^{\prime},Y^{\prime};J|Z^{\prime})$
$=I(X;Y^{\prime}|X^{\prime},J)+I(X;J|X^{\prime},Y^{\prime},Z^{\prime})$
$=I(X;Y^{\prime}|X^{\prime},J)-I(X;J|X^{\prime},Y^{\prime},Z^{\prime})$
$\leq I(Y^{\prime},Z^{\prime};X|X^{\prime})-I(J;X|X^{\prime})$
$\leq I(Y^{\prime},Z^{\prime};X|X^{\prime}).$
@@ B.6 Proof of Theorem 6
Let $F_{1:r}=(F_{1},\cdots,F_{r})$. Take some arbitrary $P_{J|X,Y,Z}$ and suppose $(X^{n},Y^{n},Z^{n},J^{n})$ are $n$ i.i.d. copies according to $P_{X,Y,Z}P_{J|X,Y,Z}$. Define the auxiliary random variables as $V_{i}=(J^{i-1},Z_{i+1}^{n},F_{1:r})$ and $U_{i}=K_{A}$. The following chain of equations follows:
@@
$\overset{(b)}{=}n\epsilon_{n}+\sum_{i=1}^{n}I(K_{A};J_{i}|J^{i-1},Z_{i+1}^{n},F_{1:r})-I(K_{A};Z_{i}|J^{i-1},Z_{i+1}^{n},F_{1:r})+H(K_{A}|J^{n},F_{1:r})$
$\overset{(c)}{=}n\epsilon_{n}+n\delta_{n}+\sum_{i=1}^{n}I(U_{i};J_{i}|V_{i})-I(U_{i};Z_{i}|V_{i})+I(K_{A};K_{B}|J^{n},F_{1:r})$
@@
where $\epsilon_{n},\delta_{n}\rightarrow 0$ as $n\rightarrow\infty$; steps (a) and (c) are derived using Fano’s inequality, (b) follows from the Marton–Körner identity, and (d) is a consequence of the Data Processing Inequality (DPI). To establish (e), we proceed as follows:
$I(X^{n};Y^{n}|J^{n})$ $\overset{(a)}{=}I(X^{n},F_{1};Y^{n}|J^{n})\overset{(b)}{\geq}I(X^{n};Y^{n}|J^{n},F_{1})\overset{(c)}{=}I(X^{n};Y^{n},F_{2}|J^{n},F_{1})\overset{(d)}{\geq}I(X^{n};Y^{n}|J^{n},F_{1},F_{2})$
@@
where (a) and (b) follow from the Markov chains $(Y^{n},F^{i-1})-\!\!\!\!\minuso\!\!\!\!-X^{n}-\!\!\!\!\minuso\!\!\!\!-F_{i}$ for odd $i$ and $(X^{n},F^{i-1})-\!\!\!\!\minuso\!\!\!\!-Y^{n}-\!\!\!\!\minuso\!\!\!\!-F_{i}$ for even $i$, respectively. Also, (b) and (d) are derived from the chain rule.
Let $Q$ be uniformly distributed over $\{1,2,\cdots,n\}$ and independent of $X,Y,Z,J,F_{1:r}$. Define the random variables $V=(V_{Q},Q)$, $U=U_{Q}$, $Z=Z_{Q}$, and $J=J_{Q}$. Under these definitions, inequality (89) simplifies to:
@@
Additionally, the auxiliary random variables $V_{i}=(J^{i-1},Z_{i+1}^{n},F_{1:r})$ and $U_{i}=K_{A}$ imply the Markov chain relationship $(U,V)-\!\!\!\!\minuso\!\!\!\!-(X,Y)-\!\!\!\!\minuso\!\!\!\!-(J,Z)$. Assuming that $n$ is sufficiently large, to prove constraint (30) for the random variables $V_{i},J_{i},Z_{i}$, and $F_{1:r}$ as defined above, we proceed as follows:
@@
$\overset{(a)}{=}\frac{1}{n}\sum_{i=1}^{n}I(F_{1:r};Z_{i}|Z_{i+1}^{n})+\frac{1}{n}\sum_{i=1}^{n}I(Z_{i+1}^{n};J_{i}|J^{i-1},F_{1:r})$
$\overset{(b)}{\leq}\frac{1}{n}I(Z^{n};F_{1:r})+\frac{1}{n}\sum_{i=1}^{n}I(J^{i-1},Z_{i+1}^{n},F_{1:r};J_{i})$
@@
where step (a) follows from the Marton–Körner identity, (b) is derived using the chain rule, and (c) utilizes the assumption
@@
The proof of the cardinality bounds is standard, and we omit it. The second part of this theorem, for the special case of $Z=f(X,Y)$, follows from Lemma 3 (after applying the appropriate change of variables). ∎
@@
Suppose that $Z=g(X)$. For every joint distribution $P_{V,X,Y}$, the maximizer of I(U;Y|V)I(U;Z|V)I(U;Y|V)-I(U;Z|V) over all $P_{U|V,X}$ is achieved when $H(X|U,V,Z)=0$.